OpenStudy (anonymous):

Among all pairs of numbers whose difference is 16, find a pair whose product is as small as possible.

jimthompson5910 (jim_thompson5910):

x-y = 16 ---> y = x-16 P = x*y P = x*(x-16) P = x^2 - 16x So your task is to minimize f(x) = x^2 - 16x

OpenStudy (anonymous):

Thanks you so much!

jimthompson5910 (jim_thompson5910):

tell me what you get when you minimize f(x)

OpenStudy (anonymous):

wouldn{t it be negative 8

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

im confused then

jimthompson5910 (jim_thompson5910):

find the vertex of y = x^2 - 16x

OpenStudy (anonymous):

In my book it gives me the answers negative 8 and negative 64 i was just wondering how

OpenStudy (anonymous):

solve it

jimthompson5910 (jim_thompson5910):

do you know how to find the vertex?

OpenStudy (anonymous):

Sorta yeah

jimthompson5910 (jim_thompson5910):

the x coordinate of the vertex is x = -b/(2a) x = -(-16)/(2*1) ... Plug in a = 1 and b = -16 x = 16/2 x = 8 So the x coordinate of the vertex is 8. This means that one of the number is 8. The other number is y = x-16 ---> y = 8-16 ---> y = -8 So the two numbers are 8 and -8. Notice they are 16 units apart (so they are separated by a distance of 16 units) and they multiply to -64

OpenStudy (anonymous):

oh, okay i understand it now!

OpenStudy (anonymous):

Can you help me on another one

jimthompson5910 (jim_thompson5910):

that's great, your book gave the answers in an odd format though (oh well)

jimthompson5910 (jim_thompson5910):

yeah sure

OpenStudy (anonymous):

You have 600 feet of fencing to enclose a rectangular plot. if you do not fence the side alone the river(600 - 2x), find the length and width of the plot that will maximize the area?whats the largest area that can be enclosed?

OpenStudy (anonymous):

And answers on the book are , length-300ft, width-150ft and maximum area 45000qt ft (if it helps out)

jimthompson5910 (jim_thompson5910):

Let x = length and y = width Draw a picture to get